Solve: 1/3x+9 - 2/x+3 = 2
Answer by stanbon(57387)

Solve: 1/3x+9 - 2/x+3 = 2
---
Multiply thru by 3(x+3)
-----
1 - 2*3 = 6(x+3)
-5 = 6x+18
6x = -23
===
c = -23/6
============
Cheers,
Stan H.
================